# Extraction of Singular Patterns from a Vector Field via Persistent Path Homology

### v2

@@ Abstract
The extraction of singular patterns is a fundamental problem in theoretical and practical domains due to the ability of such patterns to detect the intrinsic characteristics of vector fields. In this study, we propose an approach for extracting singular patterns from discrete planar vector fields. Our method involves converting the planar discrete vector field into a specialized digraph and computing its one-dimensional persistent path homology. By analyzing the persistence diagram, we can determine the location of singularity and segment a region of the singular pattern, which is referred to as a singular polygon. Moreover, the variations of singular patterns can also be analyzed. The experimental results demonstrate the effectiveness of our method in analyzing the centers and impact areas of tropical cyclones, positioning the dip poles from geomagnetic fields, and measuring variations of singular patterns between vector fields.
Index Terms:
vector field, singularity, singular pattern, digraph, persistent path homology.
I Introduction
@@
Extracting singular patterns from vector fields is an important problem in theoretical and practical domains. Singular patterns can detect the intrinsic characteristics of vector fields, such as divergence (convergence) patterns corresponding to sources (sinks) and rotational patterns corresponding to center points or spiral sources (sinks). They are commonly found in various vector field analysis problems, including vortex features of electromagnetic fields, vortex features generated by water currents, and wind field features of tropical cyclones. Identifying and extracting these singular patterns from vector fields is crucial to theoretical understanding and practical applications [1, 2].
In practical scenarios, the explicit expression of the vector field is frequently unknown, and only the field vectors at certain uniformly sampled grid points are accessible. In such instances, these vector fields are considered discrete. Fig. 1 illustrates examples of discrete vector fields.
@@
While there exists a long history of research on the extraction of singular patterns, classical numerical methods may encounter accuracy issues and lack the capability to compare differences between singular patterns. In this study, we propose a method to determine the position of singularities and segment regions of singular patterns. Moreover, we can analyze the variations of singular patterns of a time-varying vector field. Given that discrete vector fields are similar with grid-shaped digraphs, we convert the grid-sampled vector field $X$ into a specialized digraph and analyze its features using persistent path homology. Path homology [3], also known as GLMY homology, is a powerful tool for analyzing digraphs because it is sensitive to the directions in digraphs. Moreover, persistent path homology is a new perspective within the field of topological data analysis; it is specifically designed for analyzing digraph models. If a digraph contains intrinsic features conveyed by the direction and weight of its edges, these features can be discovered using (persistent) path homology.
@@
Specifically, our method involves the following steps: First, we create a special digraph called an angle-based grid digraph from $X$ and construct the corresponding digraph filtration. Then, the one-dimensional persistent path homology is computed to obtain the persistence diagram. Finally, the position of singularity can be determined, and a region of the singular pattern can be segmented, which we refer to as a singular polygon, based on the information provided by the persistence diagram. If a time-varying vector field is given, then we can also measure the changes in the topological features of the vector field by computing the distance between persistence diagrams. Moreover, in our experiments, the proposed method is employed to analyze tropical cyclones and geomagnetic fields by converting their planar vector field data into a digraph and subsequently extracting the corresponding singular patterns using persistent path homology.
Figure 1: Examples of discrete vector fields (left to right: divergence pattern, convergence pattern, and rotational pattern)
@@
  • The positions of the singularities and the corresponding regions of singular patterns are determined by computing the one-dimensional persistent path homology of digraph filtrations.
@@
  • Our method is successfully utilized to track two tropical cyclones and position the Earth’s dip poles.
The remainder of this paper is organized as follows: We provide an overview of persistent path homology and discrete vector field analysis in Section II. Then, we present a brief preliminary of vector field, digraph and path homology in Section III. The proposed method is introduced in Section IV and Section V. In Section VI we show our experiment results and provide a discussion. Section VII presents the conclusions and future works.
II Related work
@@
II-A Persistent Path Homology with Applications
@@
II-B Discrete Vector Field Analysis
@@
We now discuss methods for detecting singularities in discrete vector fields. If the expression of the vector field is known, the position of the singularity can be found by solving the system of equations [22]. However, this method is not applicable to discrete vector fields. Therefore, methods specifically designed for the discrete case have been proposed. One useful technique is Hodge decomposition [23, 24], which decomposes vector fields into three intuitive components: a divergence-free part, a curl-free part, and a harmonic part. Singularities can be detected using the decomposed vectors. Additionally, there are other numerical methods such as triangular linear interpolation of vectors [25] and the computation of the Jacobian matrix of the discrete vector field [26]. Moreover, Wong and Yip [1] introduced a method for finding the centers of circulating and spiraling vector field patterns by fitting a logarithmic spiral equation and modeling the field vectors for singular patterns corresponding to center points or spiral sources (sinks). Unlike certain numerical methods that may fail to find singularities or encounter spurious singularities caused by precision issues, thus demonstrating their limitations, this study proposes a method based on persistent path homology and an angle-based grid digraph for detecting singular patterns, relying on the singularity index.
Segmentation of discrete vector fields is important when we are interested in certain parts of a field that exhibit different patterns. Chen et al. [27] proposed a top-down approach of clustering vectors using certain similarity criteria to divide the vector field into several parts with their own features. Li et al. [2] proposed a method based on Hodge decomposition, Green function, and normalized cut to find the boundaries of regions with singular patterns. Zhang et al. [28] proposed a segmentation method that uses the accumulated geometric and physical attributes along integral curves to classify their behavior. Daniels et al. [29] introduced an interactive method for extracting vector field patterns based on the observation that although similar features in a vector field may differ spatially, they have similar neighborhood features. However, the boundary division of the singular pattern remains a significant challenge so far. In this study, we employ the theory of representative cycles of persistent path homology to segment regions of singular patterns.
@@
III Preliminaries
III-A Concepts of Vector Fields
@@
(Winding number) Suppose a simple closed curve $L$ does not contain the singularities of the vector field $X=(P,Q)$. The winding number of $X$ with respect to the curve $L$ is defined as
$j=\frac{1}{2\pi}\oint_{L}\mathrm{~{}d}\arctan\frac{Q}{P}=\frac{1}{2\pi}\oint_{L}\frac{P\mathrm{~{}d}Q-Q\mathrm{~{}d}P}{P^{2}+Q^{2}}.$
Here, the direction of the curve integral on $L$ is counterclockwise.
Definition 3.
(Index of a singularity) Suppose a simple closed curve $L$ does not contain the singularities of the vector field $X=(P,Q)$, and the region surrounded by $L$ only contains the unique singularity $S(x_{0},y_{0})$ of the vector field $X$, then the winding number of $X$ with respect to the curve $L$ is defined as the index of this singularity $S$.
Without causing confusion, the winding number of a curve is also called the index of the curve.
Theorem 1.
[22] The winding number of a simple closed curve that encloses several singularities is the sum of the respective indices of those singularities. Thus a simple closed curve that encloses no singularity has index 00.
Theorem 2.
[22] Suppose $X$ is a vector field created by piecewise linear interpolation of a discrete vector field, then the index of a singularity of $X$ is either $1$ or $-1$. If we regard $X$ as a two dimensional systems, then a singularity has index $-1$ if and only if it is a saddle.
A precise definition for singular patterns, even for basic patterns such as vortices, does not exist [31]. However, given that singular patterns are characterized by the surrounding flow field of singularity, we define the region of singular pattern in Section IV through the properties of the angle-based digraph of the field. Moreover, we define the singular polygon in Section IV to represent a region of the singular pattern.
III-B Digraph and Its Path Homology
Definition 4.
(Digraph) A digraph is an ordered pair $G=(V,E)$, where $V$ is a set of all vertices and $E$ is a set of ordered pairs of vertices, i.e. directed edges that satisfy $E\subseteq V\times V$. If $G=(V,E)$ does not contain any loop and multiple edge, it is called a simple digraph.
@@ III-B Digraph and Its Path Homology
There are five allowed 2-paths $e_{dba}$, $e_{dca}$, $e_{dgf}$, $e_{fca}$, $e_{gfc}$ that generate $\mathcal{A}_{2}$. But there is only one $\partial$-invariant 1-path $e_{dba}-e_{dca}$ that generates $\Omega_{2}$, since
@@
Now consider cycles $C_{1}=e_{db}+e_{ba}-e_{ca}-e_{dc}$ and $C_{2}=e_{dg}+e_{gf}+e_{fc}-e_{dc}$, then $C_{1},C_{2}$ are cycles that generate $Z_{1}$. But only $C_{1}$ is 1-boundary which generates $B_{1}$, since $C_{1}=\partial(e_{dba}-e_{dca})$. Hence, $H_{1}$ is generated by the class homologous to $C_{2}$ and $\dim H_{1}=1$.
@@
Figure 2: Example of persistence diagram.
@@
Figure 3: Four kinds of squares, where only the rightmost one has a trivial one-dimensional homology group. We call the rightmost one boundary square and call the others non-boundary squares.
@@
IV Extraction of singular patterns from vector fields
In this section, we develop an algorithm for extracting singularities and obtaining the corresponding singular polygons. We first introduce the angle-based grid digraph, followed by the main algorithm and the analysis of its complexity.
The process of the developed method unfolds as follows: Initially, we choose a specific area and extract the corresponding vector field data. Then we construct an angle-based grid digraph, where each edge’s weight is recorded. Subsequently, a digraph filtration is constructed by initially including all vertices of $G$ and subsequently adding edges of $G$ in a non-decreasing order of weights. This process enables the computation of the one-dimensional persistence diagram, thereby enabling the determination of the singularity’s position and the extraction of the singular polygon, as outlined in Algorithm 1. If the vector field is time-varying, the changes of singular patterns can also be identified. The overall schematic of our methods is illustrated in Fig. 8. This systematic approach ensures a comprehensive analysis of the vector field data and facilitates the identification and characterization of singular patterns.
@@ IV Extraction of singular patterns from vector fields
(A2) The singularities always have a positive index, i.e., saddles are not within our range of discussion.
@@
IV-A Constructing Angle-based Grid Digraph
@@
Figure 4: (Top) The field vector rotates clockwise when moving along edge $AB$; thus, the direction of $AB$ is $B\to A$. (Bottom) The field vector rotates counterclockwise when moving along edge $AB$; thus, the direction of $AB$ is $A\to B$.
Let $X$ be a discrete vector field defined on $D\subseteq\mathbb{R}^{2}$, sampled at uniform grid points where the distance between adjacent points in the horizontal or vertical direction is a constant $\varepsilon>0$. We define $G(X,\varepsilon)=(V,E)$ as an angle-based $\varepsilon$-grid digraph created by $X$. When the parameter $\varepsilon$ is clear from the context and does not cause confusion, we usually omit it and refer to $G(X)$ as the grid digraph in short. The vertex set $V$ consists of all the sampling points of $X$ on $D$, and $E$ is the set of directed edges determined by the following rules (Fig. 4 and 5 ): for any two adjacent vertices $A$ and $B$ in $V$ (i.e. the distance between $A$ and $B$ is exactly $\varepsilon$), if the field vector $\textbf{{v}}_{A}$ rotates counterclockwise along the line segment $AB$ to reach the field vector $\textbf{{v}}_{B}$, then the directed edge is defined as $A\to B$. Conversely, if the field vector $\textbf{{v}}_{A}$ rotates clockwise along the line segment $AB$ to reach the field vector $\textbf{{v}}_{B}$, then the directed edge is defined as $B\to A$. The weight of a directed edge $AB$ is the absolute value of the rotation angle between the two vectors $\textbf{{v}}_{A}$ and $\textbf{{v}}_{B}$. Additionally, no directed edges are found between non-adjacent vertices. Fig. 5 provides an example of an angle-based grid digraph created by a vector field.
According to assumption (A1), in the grid digraph $G$, no edge vanishing occurs between any two adjacent points. Thus, the minimal cells of $G$ consist of boundary squares and non-boundary squares. According to Theorem 4, the boundary squares in $G$ generate $B_{1}(G)$. Consequently, the minimal generators of the one-dimensional path homology group $H_{1}(G)$ can be selected as all non-boundary squares in $G$.
Figure 5: Example of angle-based grid digraph created by a vector field. The vertices of the digraph are the exact points where vectors are sampled. If the left figure is a neighborhood of the singularity, then the corresponding grid digraph satisfies theorem 5.
Now we show that transformation from vector field into digraph can retain the information of singular patterns in the following theorem. The proof of this theorem is shown in the appendix.
Theorem 5.
For example, as shown in Fig. 5, all edges of each horizontal and vertical line on the digraph have the same direction. According to the above theorem, when moving in the same direction along two horizontal or vertical lines on either side of the singularity, the direction of rotation of the field vectors are opposite. And $U$ is a region that shows the features of the singular pattern. Thus we can define the region of the singular pattern through the property of the grid digraph above.
Definition 11.
(Region of singular pattern) Suppose a vector field $X$ exists with a singularity $S$; if a neighborhood $U$ of $S$ satisfies the same condition as in Theorem 5, then $U$ is called a region of singular pattern with respect to $S$.
IV-B Extracting Position of Singularity
@@
Based on above discussions, the following theorem tells us that singularities can only exist in the interior area of non-boundary squares.
@@
Proof.
It is sufficient to prove that the singularities could not exist in the interior area of the boundary squares. Without loss of generality, let us consider the boundary square: ${C}$${D}$${A}$${B}$. According to the definition of the angle-based grid digraph, when moving from $A$ to $B$ and from $B$ to $D$, the vector undergoes a counterclockwise rotation, with the rotation angle ranging between $(0,\pi)$. Conversely, when moving from $D$ to $C$ and from $C$ to $A$, the vector rotates clockwise, with the rotation angle ranging between $(-\pi,0)$. By utilizing Equation (1), we observe that along the closed curve $ABDCA$, the sum of rotation angles falls within the range of $(-2\pi,2\pi)$. Consequently, the number of revolutions made by the vector field $X$ while traveling around the curve $ABDCA$ is zero because 0 is the only integer within the interval $(-1,1)$. As stated in Theorem 1, this implies the absence of any singularities of $X$ within the interior of the boundary square. ∎
@@
To extract the positions of singularities using one-dimensional persistent path homology, certain properties of singular patterns within a planar vector field $X$ must be considered. If a singular pattern exists in $X$, it must satisfy the following properties:
@@
The smallest square in which the singularity is located must have a shape .
Proof.
This property can be derived by Theorem 5 and Theorem 6 that the opposite edges of this square must have opposite directions. ∎
@@
The smallest square containing a singularity must have an edge with the locally greatest weight.
Proof.
We can establish this property by considering the weights of the edges in a grid digraph, which correspond to the rotation angles between the vectors at their respective endpoints. Given that all edges in a grid digraph are of the same length, as shown by Theorem 5, as we approach the singularity, the field vector on an edge undergoes a faster rotation. Consequently, the rotation angle between the vectors at both ends of an edge increases as the edge gets closer to the singularity. Therefore, the weight of an edge is greater when it is closer to the singularity. Hence, the edge with the locally greatest weight must be part of the smallest square containing the singularity. ∎
Property 3.
@@
Proof.
Before adding the edge with the greatest weight, according to Theorem 5, the representation containing the singularity must be either a or a configuration. Therefore, the edge with the greatest weight that is added last must be positioned in the middle of this hexagon. Regardless of the direction of the added edge, a new representation of the minimum generator of $H_{1}(G)$ emerges. Specifically, it is the smallest square that contains the singularity, taking the shape of a . ∎
In this property, the last square containing the singularity is formed by the edge with the greatest weight. Let the weight of this edge be denoted as $a$. This edge does not disappear in the digraph filtration, and therefore, it corresponds to a point $(a,+\infty)$ on the persistence diagram $Dgm_{1}$. Consequently, we can identify all the points on $Dgm_{1}$ that take the form $(a_{i},+\infty)$, indicating the points that persist until infinity. For each $a_{i}$ and every edge in $G(X)$ with a weight of $a_{i}$, we examine the winding number of the adjacent smallest squares associated with this edge. The square with index 1 is what we are looking for.

### v3

# Analyzing Singular Patterns in Discrete Planar Vector Fields via Persistent Path Homology

@@ Abstract
Analyzing singular patterns in vector fields is a fundamental problem in theoretical and practical domains due to the ability of such patterns to detect the intrinsic characteristics of vector fields. In this study, we propose an approach for analyzing singular patterns from discrete planar vector fields. Our method involves converting the planar discrete vector field into a specialized digraph and computing its one-dimensional persistent path homology. By analyzing the persistence diagram, we can determine the location of singularities, and the variations of singular patterns can also be analyzed. The experimental results demonstrate the effectiveness of our method in analyzing the singular patterns of noisy real-world vector fields and measuring the variations between different vector fields.
keywords:
[…]
journal: Computers & Graphics
[…]
hwlin@zju.edu.cnHongwei Lin
1 Introduction
@@
Analyzing singular patterns in vector fields is an important problem in theoretical and practical domains. Singular patterns can detect the intrinsic characteristics of vector fields, such as divergence (convergence) patterns corresponding to sources (sinks) and rotational patterns corresponding to center points or spiral sources (sinks). They are commonly found in various vector field analysis problems, including vortex features of electromagnetic fields, vortex features generated by water currents, and wind field features of tropical cyclones. Identifying and extracting these singular patterns from vector fields is crucial to theoretical understanding and practical applications [1, 2]. In practical scenarios, the explicit expression of the vector field is frequently unknown, and only the field vectors at certain uniformly sampled grid points are accessible. In such instances, these vector fields are considered discrete.
@@
While there exists a long history of research on the analysis of singular patterns, classical numerical methods may encounter accuracy issues and lack the capability to compare differences between singular patterns. In this study, we propose a new method to determine the position of singularities. Moreover, we can analyze the variations of singular patterns of a time-varying vector field. Given that discrete vector fields are similar to grid-shaped digraphs, we convert the grid-sampled vector field $X$ into a specialized digraph and analyze its features using persistent path homology. Path homology [3], also known as GLMY homology, is a powerful tool for analyzing digraphs because it is sensitive to the directions in digraphs. Moreover, persistent path homology is a new perspective within the field of computational topology; it is specifically designed for analyzing digraph models. If a digraph contains intrinsic features conveyed by the direction and weight of its edges, these features can be discovered using (persistent) path homology.
@@
Specifically, our method involves the following steps: First, we create a special digraph called an angle-based grid digraph from the vector field $X$ and construct the corresponding digraph filtration. Then, the one-dimensional persistent path homology is computed to obtain the persistence diagram. Finally, the position of singularities can be determined based on the information provided by the persistence diagram. Additionally, suppose a time-varying vector field is given. In that case, we can also measure the changes in the topological features of the vector field by computing the distance between persistence diagrams. In our experiments, the proposed method is employed to analyze geomagnetic fields and tropical cyclones by converting their real-world planar vector field data into a digraph and subsequently extracting the corresponding singular patterns using persistent path homology.
@@
  • 2.
    The positions of the singularities are determined by computing the one-dimensional persistent path homology of the digraph filtration.
  • 3.
@@
The remainder of this paper is organized as follows: We provide an overview of persistent path homology and discrete vector field analysis in Section 2. Then, we present a brief preliminary of vector field, digraph, and path homology in Section 3. The proposed method is introduced in Section 4 and Section 5. In Section 6 we show our experiment results and provide a discussion. Section 7 presents the conclusions and future works.
2 Related work
@@
2.1 Persistent path homology with applications
@@
2.2 Discrete vector field analysis
@@
We now discuss methods for detecting singularities in discrete vector fields. If the expression of the vector field is known, the position of the singularity can be found by solving the system of equations [22]. However, this method is not applicable to discrete vector fields. Therefore, methods specifically designed for the discrete case have been proposed. First, there are some numerical methods such as triangular linear interpolation of vectors [23] and the computation of the Jacobian matrix of the discrete vector field [24]. Another useful technique is Hodge decomposition [25, 26], which decomposes vector fields into three intuitive components: a divergence-free part, a curl-free part, and a harmonic part. Singularities can be detected using the decomposed vectors. Moreover, Wong and Yip [1] introduced a method for finding the centers of circulating and spiraling vector field patterns by fitting a logarithmic spiral equation and modeling the field vectors for singular patterns corresponding to center points or spiral sources (sinks). Crane et al. [27] proposed a connection defined on cells that can be used to design vector fields, and its underlying ideas can also be applied to find the locations of singularities. This method assigns vectors to each cell, and the singularities are considered to be located at the vertices of some cells. It is also possible to track singularities of time-varying vector fields. But our study focuses on detecting the accurate position of singularities within the cells. Therefore, the main purpose in [27] is different from our target. In contrast, unlike certain numerical methods that may fail to find singularities or encounter spurious singularities caused by precision issues, thus demonstrating their limitations, this study proposes a method based on persistent path homology and an angle-based grid digraph for detecting singular patterns, relying on the singularity index.
@@
3 Preliminaries
3.1 Concepts of vector fields
@@
Consider a simple closed curve $L$ that does not contain any singularities of the vector field $X=(P,Q)$. The winding number of $X$ with respect to $L$ is defined as
$j=\frac{1}{2\pi}\oint_{L}d\arctan\left(\frac{Q}{P}\right)=\frac{1}{2\pi}\oint_{L}\frac{P\,dQ-Q\,dP}{P^{2}+Q^{2}},$
where the integration along $L$ is performed counterclockwise. When the region bounded by $L$ contains exactly one singularity $S(x_{0},y_{0})$ of $X$, this winding number is called the index of $S$. For a continuous planar vector field, the winding number (or index) satisfies several fundamental properties: (1)For any simple closed curve, the index equals the sum of indices of all enclosed singularities. (2) A simple closed curve enclosing no singularities has index 00.
In this study, we focus on the vector fields that have no more than one singularity in each square (which will be introduced later), and the index of each singularity is either 1 or -1 only.
3.2 Digraph and its path homology
@@
(Digraph) A digraph is an ordered pair $G=(V,E)$, where $V$ is a set of all vertices and $E$ is a set of ordered pairs of vertices, i.e. directed edges that satisfy $E\subseteq V\times V$. If $G=(V,E)$ does not contain any self-loop, it is called a simple digraph.
For example, the digraph
${a}$${b}$${c}$
is not a simple digraph, since it has a self-loop at node $c$.
@@
It is a simple digraph, and there are five allowed 2-paths $e_{dba}$, $e_{dca}$, $e_{dgf}$, $e_{fca}$, $e_{gfc}$ that generate $\mathcal{A}_{2}$. But there is only one $\partial$-invariant 2-path $e_{dba}-e_{dca}$ that generates $\Omega_{2}$, since
@@
And other 2-paths are not elements in $\Omega_{2}$. Now consider cycles $C_{1}=e_{db}+e_{ba}-e_{ca}-e_{dc}$ and $C_{2}=e_{dg}+e_{gf}+e_{fc}-e_{dc}$, then $C_{1},C_{2}$ are cycles that generate $Z_{1}$. But only $C_{1}$ is a 1-boundary which generates $B_{1}$, since $C_{1}=\partial(e_{dba}-e_{dca})$. Hence, $H_{1}$ is generated by the class homologous to $C_{2}$ and $\dim H_{1}=1$.
@@
Fig. 1: Example of a persistence diagram.
@@
Fig. 2: Four kinds of squares, where only the rightmost one has a trivial one-dimensional homology group. We call the rightmost one boundary square and call the others non-boundary squares.
@@
Hence, the following types of digraphs are all generators of $H_{1}$:
${a}$${a}$${b}$${c}$${b}$${c}$${d}$${a}$${b}$${a}$${b}$${c}$${d}$${c}$${d}$
4 Extraction of singular patterns from vector fields
In this section, we develop an algorithm for extracting singularities. We first introduce the angle-based grid digraph, followed by the main algorithm and the analysis of its complexity. Finally, the approach for measuring variation between vector fields is introduced.
The process of the developed method unfolds as follows: Initially, we choose a specific area and extract the corresponding vector field data. Then we construct an angle-based grid digraph, where each edge’s weight is recorded. Subsequently, a digraph filtration is constructed by initially including all vertices of $G$ and subsequently adding edges of $G$ in a non-decreasing order of weights. This process enables the computation of the one-dimensional persistence diagram, thereby enabling the determination of the singularity’s position, as outlined in Algorithm 1. If the vector field is time-varying, the changes of singular patterns can also be identified. The overall schematic of our methods is illustrated in Fig. 5. This systematic approach ensures a comprehensive analysis of the vector field data and facilitates the identification and characterization of singular patterns.
@@ 4 Extraction of singular patterns from vector fields
4.1 Constructing the angle-based grid digraph
@@
Fig. 3: (Top) The field vector rotates clockwise when moving along edge $AB$; thus, the direction of $AB$ is $B\to A$. (Bottom) The field vector rotates counterclockwise when moving along edge $AB$; thus, the direction of $AB$ is $A\to B$.
Let $X$ be a discrete vector field defined on $D\subseteq\mathbb{R}^{2}$, sampled at uniform grid points where the distance between adjacent points in the horizontal or vertical direction is a constant $\varepsilon>0$. We define $G(X,\varepsilon)=(V,E)$ as an angle-based $\varepsilon$-grid digraph created by $X$. When the parameter $\varepsilon$ is clear from the context and does not cause confusion, we usually omit it and refer to $G(X)$ as the grid digraph in short. The vertex set $V$ consists of all the sampling points of $X$ on $D$, and $E$ is the set of directed edges determined by the following rules (Fig. 3 and 4 ): for any two adjacent vertices $A$ and $B$ in $V$ (i.e. the distance between $A$ and $B$ is exactly $\varepsilon$), if the field vector $\textbf{{v}}_{A}$ rotates counterclockwise along the line segment $AB$ to reach the field vector $\textbf{{v}}_{B}$, then the directed edge is defined as $A\to B$. Conversely, if the field vector $\textbf{{v}}_{A}$ rotates clockwise along the line segment $AB$ to reach the field vector $\textbf{{v}}_{B}$, then the directed edge is defined as $B\to A$. The weight of a directed edge $AB$ is the absolute value of the rotation angle between the two vectors $\textbf{{v}}_{A}$ and $\textbf{{v}}_{B}$. Additionally, no directed edges are found between non-adjacent vertices. Fig. 4 provides an example of an angle-based grid digraph created by a vector field.
According to assumption (A1), in the grid digraph $G$, no edge vanishing will occur between any two adjacent points. Thus, $G$ only consist of boundary squares and non-boundary squares, and we call these squares the minimal cells. According to Theorem 2, the boundary squares in $G$ generate $B_{1}(G)$. Consequently, the minimal generators of the one-dimensional path homology $H_{1}(G)$ can be selected as all non-boundary squares in $G$.
Fig. 4: Example of angle-based grid digraph created by a vector field. The vertices of the digraph are the exact points where vectors are sampled.
4.2 Determining the position of singularities
@@
Based on above discussions, the following theorem tells us that singularities can only exist in the interior area of non-boundary squares. The proof of this theorem is shown in Appendix A.
@@
To extract the positions of singularities using one-dimensional persistent path homology, certain properties of singular patterns within a planar vector field $X$ must be considered. If a singular pattern exists in $X$, it must satisfy the following properties. The proofs of these properties are shown in Appendix A.
@@
Every smallest square containing a singularity must have an edge $e$ with locally greatest weight, which means the weights of all edges (except $e$) of all squares containing $e$ are less than the weight of $e$.
@@
In this property, the last square containing the singularity is formed by the edge with the greatest weight. Let the weight of this edge be denoted as $a$. This edge does not disappear in the digraph filtration, and therefore, it corresponds to a point $(a,+\infty)$ on the persistence diagram $Dgm_{1}$. Consequently, we can identify all the points on $Dgm_{1}$ that take the form $(a_{i},+\infty)$, indicating the points that persist until infinity. For each $a_{i}$ and every edge in $G(X)$ with a weight of $a_{i}$, we examine the winding number of the adjacent smallest squares associated with this edge. The square with non-zero index is what we are looking for.